OpenStudy (anonymous):

4x=3y-10 and 2y=22-5x ... common solution?

OpenStudy (anonymous):

y=11-5x/2 4x=33-15x/2-10 23x=46 x=2 y=6

OpenStudy (anonymous):

thanx ur a big help